kingme44:

Matt is making cookies. He adds 3 over 4 cup of sugar to a bowl. He then removes two 1 over 8 cups of sugar. Which of the following best explains the amount of sugar left in the bowl? (1 point) Group of answer choices 3 over 8 , because 3 over 4 minus 2 times 1 over 8 equals 3 over 4 minus 2 over 8 3 over 8 , because 3 over 4 minus 2 times 1 over 8 equals 3 over 4 minus 2 over 8 3 over 8 , because 3 over 4 minus 2 times 1 over 8 equals 3 over 4 minus 3 over 8 4 over 8 , because 3 over 4 minus 2 times 1 over 8 equals 3 over 4 minus 2 over 8

jhonyy9:

so i think this wan be in this style \[\left( \frac{ 3 }{ 4 } \right) - 2\left( \frac{ 1 }{ 8 } \right) = ?\]